# Exact Gauss-Newton Optimization for Training Deep Neural Networks

### v1

@@ 4 Convergence Analysis
Let the assumptions in Lemma 4.1 hold, and assume $\lambda_{t}\equiv\lambda$ is fixed for all $t$. Let $\bar{C}$ be a positive number such that ${\mathcal{L}}_{N}(\mathbf{w}_{t})\leq\bar{C}$ holds for all $t$. Then, if the mini-batch size $b$ is such that $b\geq\lambda^{-1}((5/2)\bar{\kappa}\bar{\sigma}-\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma})>0$, the sequence $\{\mathbf{w}_{t}\}_{t=0}^{T}$, with $T\geq 1$, satisfies
$\frac{1}{T}\sum_{t=0}^{T-1}{\mathbb{E}}[\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}]\leq\frac{4b^{2}\lambda^{2}L_{1}}{\bar{\kappa}\bar{\sigma}^{2}(b\lambda+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})T}(\bar{C}-{\mathcal{L}}_{N}^{*})+\frac{5\sigma_{g}^{2}}{b}.$ (4.6)
Corollary 4.3
Suppose that the assumptions in Theorem 4.2 hold. Then, for any $\epsilon\in(0,1)$, the sequence $\{\mathbf{w}_{t}\}$ satisfies $\frac{1}{T}\sum_{t=0}^{T-1}{\mathbb{E}}[\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}]\leq\epsilon$ after at most $T=\mathcal{O}(1/\epsilon)$ iterations, provided that $\sigma_{g}^{2}=\mathcal{O}(1/T)$.
@@ 5 Experiments
[…]
@@ 5.1 Supervised Learning
For the regression task, we select three datasets: California Housing [55], Superconductivity [27], and Diamonds [74] with $20640$, $21263$, and $53940$ training samples, respectively. For classification, we use the IMDB Reviews dataset [45] containing $25000$ instances of movie reviews. Across all problems, the model $\Phi$ is a Feedforward Neural Network (FFNN) with three dense layers of $32$, $64$ and $32$ units followed by the ReLU activation function with a total of $4449$ parameters ($d=4449$). The loss function during the training is a least-squares loss (2.10) for regression and softmax cross-entropy loss (2.12) for classification. The datasets are split into training and test sets in the proportion of $90/10$ percent. Numerical features are scaled and categorical features are one-hot encoded. To measure the performance, we plot the evolution of the evaluation metric on unseen data (test set) with respect to wall time (in seconds). The evaluation metric is Root Mean Squared Error (RMSE) for regression and accuracy for classification.
Figure 1: Learning curves on the test set for SGD, Adam and EGN. The shaded area represents ±$\pm 1$ standard deviation around the mean (thick line) for $10$ seeds.
@@ 5.2 Reinforcement Learning
We demonstrate the application of EGN to reinforcement learning in two scenarios: continuous action spaces with Linear-Quadratic Regulator (LQR) and discrete action spaces using Deep Q-Network (DQN) [48].
@@
Given a discrete time-invariant linear system with continuous states and actions, and a quadratic reward function our task is to learn the optimal value function $v^{*}(s)$ and the optimal policy $\pi^{*}(s)$ such that we maximize the cumulative return. Such problems can be solved in a data-driven fashion with the policy iteration procedure [10] (outlined in Appendix D.2). It is well known that the optimal value function is quadratic and the optimal policy function is linear [28]. Consequently, we define $\Phi$ as a quadratic function of states and actions. We track the norm of the difference between the optimal LQR controller calculated analytically knowing the system matrices and the learned weights of the model.
We select two linear systems from the Compleib set of benchmarks [39]. The first system is a deterministic model of a binary distillation tower (BDT) [18], and the second one represents the linearized vertical plane dynamics of an aircraft (UAV) with noise [34]. The results are displayed in the two leftmost charts of Figure 2 and Table 2. Both EGN and SGN outperform first-order methods by a considerable margin, with EGN enjoying slightly faster convergence in both cases, while SGN achieves a marginally lower error on the stochastic LQR upon reaching convergence.
@@
Adopting the problem formulation of [48], we aim to learn the weights of a neural network that represents a $Q$-value function $q(s,a)$ that maps states $s$ and actions $a$ into scores ($Q$-values) for a discrete set of actions. Once training is complete, the optimal policy is formed by calculating the $Q$-value of each action and choosing the highest-scoring action.
We build upon CleanRL [33] framework for running RL experiments, selecting two environments: Acrobot-v1 and Freeway-v1. Acrobot-v1 is an OpenAI gym [11] environment with a $6$-dimensional state vector and a set of $3$ discrete actions where the goal is to swing the free end of the connected joints above a given height in as few steps as possible. Freeway-v1 is a MinAtar [79] environment that emulates the original Freeway Atari game. The state is represented by a $10\times 10$ image and there are $3$ discrete actions available. For Acrobot-v1 the network $\Phi$ is a FFNN with three dense layers of $32$, $64$ and $32$ units followed by the ReLU activation function with a total of $4515$ parameters. For Freeway-v1 we design a compact CNN, comprising of a convolutional layer with sixteen $3\times 3$ filters and ReLU activation, followed by flattening, a $64$-unit dense layer with ReLU, and a final dense layer outputting $Q$-values for all actions ($d=103683$).
Figure 2: Learning curves for SGD, Adam and EGN. The shaded area represents ±$\pm 1$ standard deviation around the mean return (thick line) for $10$ seeds.
@@
EGN $0$$\pm$ 0.000 $0.043\pm 0.031$81.796 $\pm$ 11.778 58.916 $\pm$ 1.946
SGN $0$$\pm$ 0.000 0.033 $\pm$ 0.019 $-114.052\pm 39.914$ $57.424\pm 3.611$
The cumulative returns from each completed episode are recorded and displayed in the rightmost charts of Figure 2. The results for both Acrobot-v1 and Freeway-v1 show no distinct advantage among the solvers, as they all reach similar episodic returns. We notice, however, that EGN slightly outperforms other optimizers by achieving a higher return level at convergence (see Table 2).
@@ 5.3 Limitations
While delivering promising results, EGN also has some limitations, as we discuss below.
@@
In our experiments we observed that the cost of computing the derivatives and the subsequent cost of computing the step are approximately the same. However, for large batch sizes ($b>128$) and parameter dimensions ($d>10^{6}$), we observed that the computational times increased significantly. We suspect this to be related to hardware limitations, and future research will aim to investigate this aspect in further detail.
@@
By incorporating the softmax function into the loss equation (2.12) we introduce coupling between the individual outputs of $\Phi$ in the denominator $\sum_{j=1}^{c}e^{\mathbf{z}_{i,j}}$, which makes the Jacobian $\mathbf{J}$ a dense $bc\times d$ matrix. Speeding up the calculation of $\mathbf{J}$ remains an open research question and constitutes a major obstacle in using Gauss-Newton methods for tasks involving a large number of classes, e.g., LLM pre-training. One possible solution could consist in moving the softmax function directly into the model $\Phi$ as the last layer of the network and then only computing the Jacobian wrt the correct class, resulting in $\mathbf{J}\in\mathbb{R}^{b\times d}$. This approach, however, yields a Hessian approximation that becomes singular close to the solution, which results in numerical instabilities that hinder convergence [25]. Another possibility suggested in [54] consists in replacing the softmax cross-entropy loss with the multi-class hinge loss. Although the computation becomes faster for the hinge loss, empirical evidence shows that the test set accuracy upon training completion is higher for the CE loss. Finally, a promising approach is the Gauss-Newton-Bartlett (GNB) estimator proposed by [42], which replaces the exact LM Hessian with an approximation obtained by sampling the subset of predicted labels.
6 Conclusions and Future Work
@@
We presented the EGN algorithm, a stochastic second-order method that efficiently estimates the descent direction by using a low-rank Gauss-Newton Hessian approximation and leveraging the Duncan-Guttman matrix identity. We demonstrated that EGN solves the system $\mathbf{H}^{\text{LM}}\mathbf{d}=-\mathbf{g}$ exactly with less computational burden than other exact Gauss-Newton methods, as well as inexact methods that rely on conjugate gradient iterates. We also proved that under mild assumptions our algorithm converges to an $\epsilon$-stationary point at a linear rate. Our empirical results show that EGN consistently matches or exceeds the generalization performance of well-tuned SGD, Adam, and SGN optimizers across various supervised and reinforcement learning tasks.
@@
Future work will focus on addressing the shortcomings of EGN in classification problems with a large number of classes. A promising direction is to approximate the Gauss-Newton Hessian matrix to avoid computing the full Jacobian of the network, e.g., using techniques such as the Gauss-Newton-Bartlett estimator [42]. Another direction is to study the performance of EGN on larger datasets and more complex models.
Appendix A Related Work
Our method can be viewed within the broader context of approximate second-order stochastic optimization. Some notable approaches include diagonal scaling [78, 42], Krylov subspace descent [73], Hessian-free optimization [46, 47, 37], quasi-Newton approach [65, 14, 5, 75, 41], Gauss-Newton  [7, 24] and Natural Gradient [3, 38] methods. A detailed overview of second-order optimization methods for large-scale machine learning problems can be found in in [8, 66, 77].
Most closely related to our work are the algorithms inspired by the Gauss-Newton approach. Those are the methods that approximate the Hessian of the loss function by utilizing first-order sensitivities only. In practice, the damped version of the Gauss-Newton direction is often calculated, forming the stochastic Levenberg-Marquardt (SLM) group of algorithms. We can classify SLM methods by three dimensions: (a) by the type of the Jacobian estimation algorithm used; (b) by the matrix inversion algorithm; and (c) by additional adaptive parameters and acceleration techniques. Based of this paradigm we summarize selected SLM algorithms in Table 3.
SGN [24] Exact via reverse mode autodiff Approximate with CG -
LM [59] Exact via reverse mode autodiff Approximate with CG Line search, momentum, uphill step acceptance
SGN [70] Exact via reverse mode autodiff Approximate with ADPG -
SGN2 [70] Approximate with SARAH estimators Approximate with ADPG -
NLLS1, NLLSL [13] Rank-1, Rank-L approximation Exact with SWM formula -
SMW-GN [60] Exact via reverse mode autodiff Exact with SWM formula Adaptive regularization
EGN (ours) Any Jacobian estimation algorithm (exact via backpropagation is the default) Exact with DG identity Line search, adaptive regularization, momentum
The Jacobian can either be calculated exactly through the reverse mode of automatic differentiation as proposed by, e.g., [24, 13, 70] or be estimated approximately. Low rank Jacobian estimation is suggested by the NLLS1 and NLLSL algorithms [13] with experimental results showing almost on par performance with the exact Jacobian version of the methods. SGN2 [70] uses SARAH estimators for approximating function values and Jacobians with SGN2 performing better than SGN [70] that assumes the exact Jacobian. In [42] the Gauss-Newton-Bartlett (GNB) estimator is introduced to adapt the GN method to large-scale classification problems. EGN does not mandate a specific computation technique for the Jacobian. In our experiments we rely on backpropagation deferring other methods to further research.
Solving $\mathbf{H}\mathbf{d}=-\mathbf{g}$ naively for a neural network with $d$ parameters has complexity $\mathcal{O}\left(d^{3}\right)$. The procedure becomes practically infeasible even for networks of moderate size, so several alternative approaches have been proposed. Following [6], we distinguish between inexact and exact solutions to the linear system. Inexact methods rely on iterative algorithms to solve the system approximately in as few iterations as possible. Among such algorithms we mention the Conjugate Gradient (CG) method used in Hessian-free optimization [47, 37] as well as in GN methods like SGN [24] and LM [59]; the Accelerated Dual Proximal-Gradient (ADPG) method proposed in [70] for SGN and SGN2 solvers; the Stochastic Gradient Iteration approach (SGI), analysed in [6] as part of the Newton-SGI solver. Exact methods are typically based on linear algebra identities. For example, in [60] and [13] the system is solved exactly with the Sherman-Morrison-Woodbury (SMW) formula. Contrarily, we follow  [1] and derive the EGN update formula using the Duncan-Guttman matrix identity [23, 26].
State-of-the-art implementations of the GN algorithm often include additional improvements to address the issues of stochasticity of the Hessian matrix, computational load of calculating the Jacobian and adaptive hyper-parameters tuning. Just as with the gradient, the stochastic sampling introduces noise in the Hessian which leads to the erroneous descent direction. A common solution to combat noisy estimates is to add temporal averaging (or momentum), eg. with exponential moving averages (EMA). Examples of such approach are AdaHessian [78] and Sophia [42] that keep EMA of a diagonal Hessian, as well as  [37] that incorporates momentum into the HFO framework. The idea of re-using the Hessian estimate from previous iterations is formalized in [20] showing that evaluating the Hessian “lazily” once per $k$ iterations significantly reduces the computational burden while at the same time does not degrade the performance that much. Although second-order methods typically require less tuning [7, 24, 77], the SLM approach still requires setting a learning rate $\alpha$ and the regularization parameter $\lambda$. The line search for $\alpha$, common in the deterministic optimization, is problematic in the stochastic setting due to the high variance of the loss gradient norm [17]. Still, there are promising attempts to incorporate line search into quasi-Newton methods [75], HFO [37] as well as Gauss-Newton [59]. Adaptive regularization in a manner similar to the deterministic Levenberg-Marquardt approach is proposed in [31, 60].
Appendix B Proofs and Derivations
B.1 Generalized Gauss-Newton Hessian Approximation

### v2

@@ 5 Convergence Analysis
Let the assumptions in Lemma 5.5 hold, and assume $\lambda_{t}\equiv\lambda$ is fixed for all $t$. Let $\alpha_{t}=\frac{\alpha_{0}}{(t+1)^{a}}$ for some $0<\alpha_{0}<1$, $\frac{1}{2}<a<1$. Then, the loss gradient approaches 0 in expectation as the iteration count $T\to\infty$.
Proof. Following the proof of Lemma 5.5, we fix $\alpha^{\text{max}}=1/c_{1}$ in Algorithm 4, with $c_{1}=(5b^{2}\lambda_{t}L_{1})/(3(b\lambda_{t}+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})^{2}$, and take the conditional expectation on both sides of (5.6). This yields
${\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t+1})]$ $\leq{\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t})]-\frac{1}{2}\alpha_{t}{\mathbb{E}}[\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}]+C\alpha_{t}^{2},$ (5.12)
where $C:=\frac{3\sigma_{g}^{2}c_{1}}{10b\lambda_{t}}$. Summing (5.12) over $t=0,1,\ldots,T-1$, we obtain
${\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{T})]$ $\leq{\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{0})]-\frac{1}{2}\sum_{t=0}^{T-1}\alpha_{t}{\mathbb{E}}[\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}]+C\sum_{t=0}^{T-1}\alpha_{t}^{2}.$ (5.13)
Assuming that the learning rates are given by $\alpha_{t}=\frac{\alpha_{0}}{(t+1)^{a}}$ for some $0<\alpha_{0}<1$, $\frac{1}{2}<a<1$, we have
$\sum_{t=0}^{T-1}\alpha_{t}^{2}=D<\infty.$
Next, consider a random variable $\mathbf{z}_{T}$ satisfying $P[\mathbf{z}_{T}=\mathbf{w}_{t}]=\frac{1}{T},$ for all $t$. Consequently,
${\mathbb{E}}[\|\nabla{\mathcal{L}}_{N}(\mathbf{z}_{T})\|^{2}]=\frac{1}{T}\sum_{t=0}^{T-1}{\mathbb{E}}[\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}],$
such that, using also $\alpha_{t_{2}}\leq\alpha_{t_{1}}$ for all $t_{2}\geq t_{1}$, (5.13) becomes
${\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{T})]$ $\leq{\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{0})]-\alpha_{T-1}\frac{1}{2}T{\mathbb{E}}[\|\nabla{\mathcal{L}}_{N}(\mathbf{z}_{T})\|^{2}]+CD.$
Then, using ${\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{T})]\geq{\mathcal{L}}_{N}^{\star}$ we have
${\mathbb{E}}[\|\nabla{\mathcal{L}}_{N}(\mathbf{z}_{T})\|^{2}]$ $\leq 2\frac{{\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{0})]-{\mathcal{L}}_{N}^{*}}{\alpha_{T-1}T}+\frac{2CD}{\alpha_{T-1}T}=2\frac{{\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{0})]-{\mathcal{L}}_{N}^{*}}{\alpha_{0}T^{1-a}}+\frac{2CD}{\alpha_{0}T^{1-a}}.$
Taking the limit for $T\to\infty$ we finally have
$\lim_{T\to\infty}{\mathbb{E}}[\|\nabla{\mathcal{L}}_{N}(\mathbf{z}_{T})\|^{2}]$ $\leq 0.$
$\Box$
@@ 6 Experiments
We conduct a series of experiments to measure the performance of EGN on several supervised learning and reinforcement learning tasks.
We select five baseline solvers: SGD robbins1951stochastic , Adam kingma2014adam , SGD with momentum and Gradient Activation Function (GAF) liu2021activated , a Quasi-Newton solver (SQN) bottou2018optimization ; byrd2016stochastic ; schraudolph2007stochastic , and SGN gargiani2020promise . SGD acts as a most basic baseline with the computationally cheapest update; Adam is a widely used accelerated FOM for training DNNs; GAF is a recent first-order variant that reports faster convergence on deep networks; SGN is an inexact Gauss-Newton solver against which we evaluate the practical advantages of solving the system (3.5) exactly; and SQN acts as an alternative to Gauss-Newton that approximates the Hessian via low-rank updates. Since first-order methods typically require fewer computations per iterate, in order to obtain a fair comparison we monitor the wall time instead of the number of iterations. The learning rates are selected as the best performing $\alpha$ after a grid search in the logspace $\alpha\in[10^{-9},1]$. Additionally, for SGN we search for the optimal “number of CG iterations” within the set $\{3,5,10,20,50\}$. For EGN we introduce two extra hyper-parameters: “line search” $\{\text{True},\text{False}\}$ and “momentum” $\{0.0,0.9\}$. The best performing sets of hyper-parameters as well as detailed description of the datasets are available in Appendix C. The size of the mini-batch for all problems is $128$. All the experiments are conducted on the Tesla T4 GPU in the Google Colab environment with float32 precision.
@@ 6.1 Supervised Learning
For the regression task, we select three datasets: California Housing ds_california_housing , Superconductivity ds_superconductivty , and Diamonds ds_diamonds with $20640$, $21263$, and $53940$ training samples, respectively. For classification, we use the IMDB Reviews dataset ds_imdb containing $25000$ instances of movie reviews. Across all problems, the model $\Phi$ is a Feedforward Neural Network (FFNN) with three dense layers of $32$, $64$ and $32$ units followed by the ReLU activation function with a total of $4449$ parameters ($d=4449$). The loss function during the training is a least-squares loss (3.10) for regression and softmax cross-entropy loss (3.12) for classification. The datasets are split into training and test sets in the proportion of $90/10$ percent. Numerical features are scaled and categorical features are one-hot encoded. To measure the performance, we plot the evolution of the evaluation metric on unseen data (test set) with respect to wall time (in seconds). The evaluation metric is Root Mean Squared Error (RMSE) for regression and accuracy for classification.
Figure 1: Learning curves on the test set for SGD, Adam, GAF, SQN, SGN and EGN. The shaded area represents $\pm 1$ standard deviation around the mean (thick line) for $10$ seeds.
@@ 6.2 Reinforcement Learning
We demonstrate the application of EGN to reinforcement learning in two scenarios: continuous action spaces with Linear-Quadratic Regulator (LQR) and discrete action spaces using Deep Q-Network (DQN) mnih2013playing .
@@
Given a discrete time-invariant linear system with continuous states and actions, and a quadratic reward function our task is to learn the optimal value function $v^{*}(s)$ and the optimal policy $\pi^{*}(s)$ such that we maximize the cumulative return. Such problems can be solved in a data-driven fashion with the policy iteration procedure bradtke1994adaptive (outlined in Appendix C.2). It is well known that the optimal value function is quadratic and the optimal policy function is linear hazan2022introduction . Consequently, we define $\Phi$ as a quadratic function of states and actions. We track the norm of the difference between the optimal LQR controller calculated analytically knowing the system matrices and the learned weights of the model.
We select two linear systems from the Compleib set of benchmarks leibfritz2006compleib . The first system is a deterministic model of a binary distillation tower (BDT) davison1990benchmark , and the second one represents the linearized vertical plane dynamics of an aircraft (UAV) with noise hung1982multivariable . The results are displayed in the top two charts of Figure 2 and Table 4. Both EGN and SGN outperform first-order methods by a considerable margin, with EGN enjoying slightly faster convergence in both cases, while SGN achieves a marginally lower error on the stochastic LQR upon reaching convergence.
@@
Adopting the problem formulation of mnih2013playing , we aim to learn the weights of a neural network that represent a $Q$-value function $q(s,a)$ that maps states $s$ and actions $a$ into scores ($Q$-values) for a discrete set of actions. Once training is complete, the optimal policy is formed by calculating the $Q$-value of each action and choosing the highest-scoring action.
Figure 2: Learning curves for SGD, Adam, SGN and EGN. The shaded area represents $\pm 1$ standard deviation around the mean return (thick line) for $10$ seeds.
@@
EGN 0.000 $\pm$ 0.000 0.043 $\pm$ 0.031 -81.796 $\pm$ 11.778 58.916 $\pm$ 1.946
SGN 0.000 $\pm$ 0.000 0.033 $\pm$ 0.019 -114.052 $\pm$ 39.914 57.424 $\pm$ 3.611
We build upon CleanRL huang2021cleanrl framework for running RL experiments, selecting two environments: Acrobot-v1 and Freeway-v1. Acrobot-v1 is an OpenAI gym openai_gym environment with a $6$-dimensional state vector and a set of $3$ discrete actions where the goal is to swing the free end of the connected joints above a given height in as few steps as possible. Freeway-v1 is a MinAtar young19minatar environment that emulates the original Freeway Atari game. The state is represented by a $10\times 10$ image and there are $3$ discrete actions available. For Acrobot-v1 the network $\Phi$ is a FFNN with three dense layers of $32$, $64$ and $32$ units followed by the ReLU activation function with a total of $4515$ parameters. For Freeway-v1 we design a compact CNN, comprising of a convolutional layer with sixteen $3\times 3$ filters and ReLU activation, followed by flattening, a $64$-unit dense layer with ReLU, and a final dense layer outputting $Q$-values for all actions ($d=103683$).
The cumulative returns from each completed episode are recorded and displayed in the bottom two charts of Figure 2. The results for both Acrobot-v1 and Freeway-v1 show no distinct advantage among the solvers, as they all reach similar episodic returns. We notice, however, that EGN slightly outperforms other optimizers by achieving a higher return level at convergence (see Table 4).
@@ 6.3 Limitations
In our experiments we observed that the cost of computing the derivatives and the subsequent cost of computing the step (Algorithm 1) are comparable. However, for large batch sizes ($b>128$) we observed that the computational times increased significantly. Figure 3 quantifies how the direction calculation stage begins to dominate the update time as the batch size grows. While EGN remains an efficient drop-in replacement for first-order methods within the commonly-used range $32\leq b\leq 128$, scaling to very large batches will require additional techniques. We suspect this to be related to hardware limitations and leave a systematic study of such mitigations to future work.
Figure 3: Proportion of total update time spent in Algorithm 1 as a function of batch size for fully-connected neural networks (MLPs) of various sizes. Curves show the mean over 1000 runs; shaded regions denote $\pm 1$ standard deviation. Absolute wall-clock times are reported in Appendix C.4.
@@
By incorporating the softmax function into the loss equation (3.12) we introduce coupling between the individual outputs of $\Phi$ in the denominator $\sum_{j=1}^{c}e^{\mathbf{z}_{i,j}}$, which makes the Jacobian $\mathbf{J}$ a dense $bc\times d$ matrix. Speeding up the calculation of $\mathbf{J}$ remains an open research question and constitutes a major obstacle in using Gauss-Newton methods for tasks involving a large number of classes, e.g., LLM pre-training. One possible solution could consist in moving the softmax function directly into the model $\Phi$ as the last layer of the network and then only computing the Jacobian with respect to the correct class, resulting in $\mathbf{J}\in\mathbb{R}^{b\times d}$. This approach, however, yields a Hessian approximation that becomes singular close to the solution, which results in numerical instabilities that hinder convergence grosse2021taylor . Another possibility suggested in ozyildirim2021levenberg consists of replacing the softmax cross-entropy loss with the multi-class hinge loss. Although the computation becomes faster for the hinge loss, empirical evidence shows that the test set accuracy upon training completion is higher for the CE loss. Finally, a promising approach is the Gauss-Newton-Bartlett (GNB) estimator proposed by liu2023sophia , which replaces the exact LM Hessian with an approximation obtained by sampling the subset of predicted labels.
7 Conclusion
@@
We presented the EGN algorithm, a stochastic second-order method that efficiently estimates the descent direction by using a low-rank Gauss-Newton Hessian approximation and leveraging the Duncan-Guttman matrix identity. We demonstrated that EGN solves the system $\mathbf{H}^{\text{LM}}\mathbf{d}=-\mathbf{g}$ exactly with less computational burden than other exact Gauss-Newton methods, as well as inexact methods that rely on conjugate gradient iterates. We also proved that under mild assumptions our algorithm converges in expectation. Our empirical results show that EGN consistently matches or exceeds the generalization performance of well-tuned SGD, Adam, GAF, SQN, and SGN optimizers across various supervised and reinforcement learning tasks.
@@
Future work will focus on addressing the shortcomings of EGN in classification problems with a large number of classes. A promising direction is to approximate the Gauss-Newton Hessian matrix to avoid computing the full Jacobian of the network, e.g., using techniques such as the Gauss-Newton-Bartlett estimator liu2023sophia . Another direction is to study the performance of EGN on larger datasets and more complex models.
Acknowledgments
This work was partially funded by the European Union (ERC Advanced Research Grant COMPACT, No. $101141351$). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Proofs and Derivations
A.1 Generalized Gauss-Newton Hessian Approximation
Claim